# Model-Agnostic Bounds for Augmented Inverse Probability Weighted Estimators’ Wald-Confidence Interval Coverage in Randomized Controlled Trials

### v1

@@ 6 Discussion
${\mathbb{E}}[\{P_{*}({\mathcal{T}}_{a}(\hat{Q})^{2}-{\mathbb{E}}_{\hat{Q}}[{\mathcal{T}}_{a}(\hat{Q})^{2}])\}^{2}]\leq 16M^{2}P_{*}\left(\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}{\mathbb{E}}_{\hat{Q}}\left[\left(\hat{Q}-Q_{\#,a}\right)^{2}\right]\right).$
@@
$16M^{2}\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}{\mathbb{E}}\|\hat{Q}-Q_{\#,a}\|_{2}^{2}.$
@@
The first four equations follow from direct calculation. Moreover, ${\mathbb{E}}[P_{*}({\mathcal{T}}_{a}(\hat{Q})^{2}-{\mathbb{E}}_{\hat{Q}}[{\mathcal{T}}_{a}(\hat{Q})^{2}])]=0$ and
${\mathbb{E}}[\{P_{*}({\mathcal{T}}_{a}(\hat{Q})^{2}-{\mathbb{E}}_{\hat{Q}}[{\mathcal{T}}_{a}(\hat{Q})^{2}])\}^{2}]$
$\leq{\mathbb{E}}\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}\left\{-2Q_{*,a}(\hat{Q}-{\mathbb{E}}_{\hat{Q}}[\hat{Q}])+\hat{Q}^{2}-{\mathbb{E}}_{\hat{Q}}[\hat{Q}^{2}]\right\}^{2}\right].$
@@
${\mathbb{E}}\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}\left\{-2Q_{*,a}(\hat{Q}-{\mathbb{E}}_{\hat{Q}}[\hat{Q}])\right\}^{2}\right]$
$=4P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}Q_{*,a}^{2}{\mathbb{E}}_{\hat{Q}}\left[(\hat{Q}-{\mathbb{E}}_{\hat{Q}}[\hat{Q}])^{2}\right]$
$\leq 4M^{2}P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}{\mathbb{E}}_{\hat{Q}}\left[(\hat{Q}-Q_{\#,a})^{2}\right],$
${\mathbb{E}}\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}\left\{\hat{Q}^{2}-{\mathbb{E}}_{\hat{Q}}[\hat{Q}^{2}]\right\}^{2}\right]$
$=P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}{\mathbb{E}}_{\hat{Q}}\left[\left(\hat{Q}^{2}-{\mathbb{E}}_{\hat{Q}}[\hat{Q}^{2}]\right)^{2}\right]$
$\leq P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}{\mathbb{E}}_{\hat{Q}}\left[\left(\hat{Q}^{2}-Q_{\#,a}^{2}\right)^{2}\right]$
$\leq 4M^{2}P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}{\mathbb{E}}_{\hat{Q}}\left[\left(\hat{Q}-Q_{\#,a}\right)^{2}\right].$
@@
${\mathbb{E}}\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}\left\{-2Q_{*,a}(\hat{Q}-{\mathbb{E}}_{\hat{Q}}[\hat{Q}])+\hat{Q}^{2}-{\mathbb{E}}_{\hat{Q}}[\hat{Q}^{2}]\right\}^{2}\right]$
$\leq\Bigg\{\sqrt{{\mathbb{E}}\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}\left\{-2Q_{*,a}(\hat{Q}-{\mathbb{E}}_{\hat{Q}}[\hat{Q}])\right\}^{2}\right]}$
$\qquad+\sqrt{{\mathbb{E}}\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}\left\{\hat{Q}^{2}-{\mathbb{E}}_{\hat{Q}}[\hat{Q}^{2}]\right\}^{2}\right]}\Bigg\}^{2}$
$\leq 16M^{2}P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}{\mathbb{E}}_{\hat{Q}}\left[\left(\hat{Q}-Q_{\#,a}\right)^{2}\right].$
The bound on ${\mathbb{E}}[\{P_{*}({\mathcal{T}}_{a}(\hat{Q})^{2}-{\mathbb{E}}_{\hat{Q}}[{\mathcal{T}}_{a}(\hat{Q})^{2}])\}^{2}]$ under Condition 3 follows immediately by upper-bounding $\{1-\pi_{*}(a\mid x)\}/\pi_{*}(a\mid x)$. ∎
S4.2 Results for cross-fit AIPW procedure
The next lemma is a tail bound for $P_{n,k}H_{a}(\hat{Q}_{k,a})$.
@@
$\Pr(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta)\leq\frac{|I_{k}|}{n}\frac{1-\tau_{\pi}}{\tau_{\pi}\eta^{2}}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}.$
Moreover, under Condition 7, for any constant $q>0$,
$\Pr(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta)$
$\leq 2\exp\left\{-\frac{n\eta^{2}}{2|I_{k}|\hat{s}_{1}^{2}\left[{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}+\hat{s}_{2}\sqrt{2q{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\log n}\right]^{2}}\right\}+2n^{-q}.$
@@
Note that $P_{n,k}H_{a}(\hat{Q}_{k,a})$ is a sample average of $H_{a}(\hat{Q}_{k,a})(V_{i})$ ($i\in I_{k}$), which are i.i.d. random variables conditional on $\hat{Q}_{k,a}$. By Lemma S5, ${\mathbb{E}}[P_{n,k}H_{a}(\hat{Q}_{k,a})\mid\hat{Q}_{k,a}]=0$ and ${\mathrm{Var}}(P_{n,k}H_{a}(\hat{Q}_{k,a})\mid\hat{Q}_{k,a})=P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}/|I_{k}|$.
First show the bound without Condition 7. By Chebyshev’s inequality,
$\Pr(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta)$ $={\mathbb{E}}\left[\Pr(|P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta/|I_{k}|\mid\hat{Q}_{k,a})\right]$
$\leq{\mathbb{E}}\left[\frac{|I_{k}|}{n\eta^{2}}P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}\right]$
$\leq\frac{|I_{k}|}{n}\frac{1-\tau_{\pi}}{\tau_{\pi}\eta^{2}}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2},$
@@
Next show the bound under Condition 7. By Hoeffding’s bound for sums of independent subgaussian random variables (e.g., Proposition 2.5 in Wainwright,, 2019),
$\Pr(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta)$ $={\mathbb{E}}\left[\Pr(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta\mid\hat{Q}_{k,a})\right]$
$\leq 2{\mathbb{E}}\left[\exp\left\{-\frac{n\eta^{2}}{2|I_{k}|\hat{s}_{1}^{2}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\right\}\right].$ (S6)
@@
$2{\mathbb{E}}\left[\exp\left\{-\frac{n\eta^{2}}{2|I_{k}|\hat{s}_{1}^{2}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\right\}{\mathbbm{1}}(\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}\leq{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}+t)\right]$
$+2{\mathbb{E}}\left[\exp\left\{-\frac{n\eta^{2}}{2|I_{k}|\hat{s}_{1}^{2}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\right\}{\mathbbm{1}}(\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}>{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}+t)\right]$
$\leq 2\exp\left\{-\frac{n\eta^{2}}{2|I_{k}|\hat{s}_{1}^{2}\left[{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}+t\right]^{2}}\right\}+2\Pr(\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}>{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}+t)$
$\leq 2\exp\left\{-\frac{n\eta^{2}}{2|I_{k}|\hat{s}_{1}^{2}\left[{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}+t\right]^{2}}\right\}+2\exp\left\{-\frac{t^{2}}{2\hat{s}_{2}^{2}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\right\}.$
Taking $t=\hat{s}_{2}\sqrt{2q{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\log n}$ and combining the above results, the desired inequality follows. ∎
@@
Let $k\in\{1,\ldots,K\}$ be arbitrary. It holds that ${\mathbb{E}}[\hat{\psi}_{k,a}]={\mathbb{E}}[\hat{\psi}_{a}]=\psi_{*,a}$,
${\mathrm{Var}}(\hat{\psi}_{k,a})$ $\leq\frac{1}{|I_{k}|}\left\{\sigma_{\#,a}+\sqrt{{\mathbb{E}}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}\right]}\right\}^{2},$
${\mathrm{Var}}(\hat{\psi}_{a})$ $\leq\left\{\frac{\sigma_{\#,a}}{\sqrt{n}}+\sum_{k=1}^{K}\frac{|I_{k}|}{n}\sqrt{{\mathbb{E}}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\frac{(\hat{Q}_{k,a}-Q_{\#,a})^{2}}{|I_{k}|}\right]}\right\}^{2}.$
@@
${\mathrm{Var}}(\hat{\psi}_{k,a})$ $\leq\frac{1}{|I_{k}|}\left\{\sigma_{\#,a}+\sqrt{\frac{1-\tau_{\pi}}{\tau_{\pi}}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\right\}^{2},$
${\mathrm{Var}}(\hat{\psi}_{a})$ $\leq\left\{\frac{\sigma_{\#,a}}{\sqrt{n}}+\sum_{k=1}^{K}\frac{\sqrt{|I_{k}|}}{n}\sqrt{\frac{1-\tau_{\pi}}{\tau_{\pi}}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\right\}^{2}.$
@@
Since $\hat{Q}_{k,a}$ and $P_{n,k}$ are based on independent data, by Lemma S5, ${\mathbb{E}}[(P_{n,k}-P_{*})H_{a}(\hat{Q}_{k,a})]=0$, and so ${\mathbb{E}}[\hat{\psi}_{k,a}]={\mathbb{E}}[\hat{\psi}_{a}]=\psi_{*,a}$.
@@
${\mathrm{Var}}((P_{n,k}-P_{*})H_{a}(\hat{Q}_{k,a}))={\mathbb{E}}[P_{*}H_{a}(\hat{Q}_{k,a})^{2}]/|I_{k}|=\frac{1}{|I_{k}|}{\mathbb{E}}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}\right].$
@@
${\mathrm{Var}}(\hat{\psi}_{k,a})={\mathbb{E}}[(\hat{\psi}_{k,a}-\psi_{*,a})^{2}]$ $\leq\left\{\sqrt{{\mathrm{Var}}[(P_{n,k}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})]}+\sqrt{{\mathrm{Var}}[(P_{n,k}-P_{*})H_{a}(\hat{Q}_{k,a})]}\right\}^{2}$
$=\frac{1}{|I_{k}|}\left\{\sigma_{\#,a}+\sqrt{{\mathbb{E}}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}\right]}\right\}^{2},$
${\mathrm{Var}}(\hat{\psi}_{a})={\mathbb{E}}[(\hat{\psi}_{a}-\psi_{*,a})^{2}]$ $\leq\left\{\sqrt{{\mathrm{Var}}((P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a}))}+\sum_{k=1}^{K}\sqrt{{\mathrm{Var}}\left(\frac{|I_{k}|}{n}(P_{n,k}-P_{*})H_{a}(\hat{Q}_{k,a})\right)}\right\}^{2}$
$=\left\{\frac{\sigma_{\#,a}}{\sqrt{n}}+\sum_{k=1}^{K}\frac{|I_{k}|}{n}\sqrt{{\mathbb{E}}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\frac{(\hat{Q}_{k,a}-Q_{\#,a})^{2}}{|I_{k}|}\right]}\right\}^{2}.$
@@
$\Pr\left(|\hat{\sigma}_{k,a}^{2}-{\mathbb{E}}[\hat{\sigma}_{k,a}^{2}]|>\frac{n}{|I_{k}|}\sum_{j=1}^{3}\lambda_{j}\right)$
@@
$\bar{\hat{\mathcal{R}}}(\lambda_{2}):=16M^{2}\frac{|I_{k}|^{2}}{n^{2}\lambda_{2}^{2}}\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}.$
Under Condition S1, $\bar{\hat{\mathcal{R}}}(\lambda_{2})$ may be replaced by
$\bar{\hat{\mathcal{R}}}^{\prime}(\lambda_{2}):=2\exp\left\{-\frac{n^{2}\tau_{\pi}^{2}\lambda_{2}^{2}}{2|I_{k}|^{2}\hat{s}_{3}^{2}M^{2}(1-\tau_{\pi})^{2}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\right\}.$

### v2

@@ 6 Discussion
${E}[\{P_{*}({\mathcal{T}}_{a}(\hat{Q})^{2}-{E}_{\hat{Q}}[{\mathcal{T}}_{a}(\hat{Q})^{2}])\}^{2}]\leq 16M^{2}P_{*}\left(\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}{E}_{\hat{Q}}\left[\left(\hat{Q}-Q_{\#,a}\right)^{2}\right]\right).$
@@
$16M^{2}\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}{E}\|\hat{Q}-Q_{\#,a}\|_{2}^{2}.$
@@
The first five equations follow from direct calculation. Moreover, ${E}[P_{*}({\mathcal{T}}_{a}(\hat{Q})^{2}-{E}_{\hat{Q}}[{\mathcal{T}}_{a}(\hat{Q})^{2}])]=0$ and
${E}[\{P_{*}({\mathcal{T}}_{a}(\hat{Q})^{2}-{E}_{\hat{Q}}[{\mathcal{T}}_{a}(\hat{Q})^{2}])\}^{2}]$
$\leq{E}\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}\left\{-2Q_{*,a}(\hat{Q}-{E}_{\hat{Q}}[\hat{Q}])+\hat{Q}^{2}-{E}_{\hat{Q}}[\hat{Q}^{2}]\right\}^{2}\right].$
@@
${E}\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}\left\{-2Q_{*,a}(\hat{Q}-{E}_{\hat{Q}}[\hat{Q}])\right\}^{2}\right]$
$=4P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}Q_{*,a}^{2}{E}_{\hat{Q}}\left[(\hat{Q}-{E}_{\hat{Q}}[\hat{Q}])^{2}\right]$
$\leq 4M^{2}P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}{E}_{\hat{Q}}\left[(\hat{Q}-Q_{\#,a})^{2}\right],$
${E}\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}\left\{\hat{Q}^{2}-{E}_{\hat{Q}}[\hat{Q}^{2}]\right\}^{2}\right]$
$=P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}{E}_{\hat{Q}}\left[\left(\hat{Q}^{2}-{E}_{\hat{Q}}[\hat{Q}^{2}]\right)^{2}\right]$
$\leq P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}{E}_{\hat{Q}}\left[\left(\hat{Q}^{2}-Q_{\#,a}^{2}\right)^{2}\right]$
$\leq 4M^{2}P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}{E}_{\hat{Q}}\left[\left(\hat{Q}-Q_{\#,a}\right)^{2}\right].$
@@
${E}\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}\left\{-2Q_{*,a}(\hat{Q}-{E}_{\hat{Q}}[\hat{Q}])+\hat{Q}^{2}-{E}_{\hat{Q}}[\hat{Q}^{2}]\right\}^{2}\right]$
$\leq\Bigg\{\sqrt{{E}\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}\left\{-2Q_{*,a}(\hat{Q}-{E}_{\hat{Q}}[\hat{Q}])\right\}^{2}\right]}$
$\qquad+\sqrt{{E}\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}\left\{\hat{Q}^{2}-{E}_{\hat{Q}}[\hat{Q}^{2}]\right\}^{2}\right]}\Bigg\}^{2}$
$\leq 16M^{2}P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\right\}^{2}{E}_{\hat{Q}}\left[\left(\hat{Q}-Q_{\#,a}\right)^{2}\right].$
The bound on ${E}[\{P_{*}({\mathcal{T}}_{a}(\hat{Q})^{2}-{E}_{\hat{Q}}[{\mathcal{T}}_{a}(\hat{Q})^{2}])\}^{2}]$ under Condition 3 follows immediately by upper-bounding $\{1-\pi_{*}(a\mid x)\}/\pi_{*}(a\mid x)$. ∎
S5.2 Results for cross-fit AIPW procedure
The next lemma is a tail bound for $P_{n,k}H_{a}(\hat{Q}_{k,a})$. The tail bound without the approximate sub-Weibull condition 7 can be improved under Condition 4, but it should not affect the rate in Theorem 1.
@@
${\mathrm{pr}}(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta)\leq\frac{|I_{k}|}{n}\frac{1-\tau_{\pi}}{\tau_{\pi}\eta^{2}}{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}.$
Moreover, under Conditions 4 and 7, for any constant $q>0$,
${\mathrm{pr}}(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta)\leq 2\exp(-\eta^{\prime})+2\hat{a}_{1}n^{-q}+2\hat{b}_{1}\sqrt{\log n/n},$
where
$\eta^{\prime}:=\eta^{\prime}(\eta):=\frac{\tau_{\pi}}{1-\tau_{\pi}}\left\{\frac{n[(q+\hat{d}_{1})\hat{c}_{1}^{-1}\log n]^{-2/\hat{q}_{1}}\eta^{2}}{4|I_{k}|\hat{r}_{1}(n)^{2}}\wedge\frac{3\sqrt{n}\eta}{8M}\right\}.$
@@
Note that $P_{n,k}H_{a}(\hat{Q}_{k,a})$ is a sample average of $H_{a}(\hat{Q}_{k,a})(V_{i})$ ($i\in I_{k}$), which are i.i.d. random variables conditional on $\hat{Q}_{k,a}$. By Lemma S5, ${E}[P_{n,k}H_{a}(\hat{Q}_{k,a})\mid\hat{Q}_{k,a}]=0$ and ${\mathrm{var}}(|I_{k}|P_{n,k}H_{a}(\hat{Q}_{k,a})\mid\hat{Q}_{k,a})=|I_{k}|P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}$.
First show the bound without Conditions 4 and 7. By Chebyshev’s inequality,
${\mathrm{pr}}(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta)$ $={E}\left[{\mathrm{pr}}(|P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta/|I_{k}|\mid\hat{Q}_{k,a})\right]$
$\leq{E}\left[\frac{|I_{k}|}{n\eta^{2}}P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}\right]$
$\leq\frac{|I_{k}|}{n}\frac{1-\tau_{\pi}}{\tau_{\pi}\eta^{2}}{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2},$
@@
Next show the bound under Conditions 4 and 7. For any $x\geq 0$, let
$\hat{\eta}(x):=\frac{\tau_{\pi}}{1-\tau_{\pi}}\left(\frac{n\eta^{2}}{4|I_{k}|x^{2}}\wedge\frac{3\sqrt{n}\eta}{8M}\right),$
where $1/0:=\infty$. Under Conditions 3 and 4, $|H_{a}(\hat{Q}_{k,a})(V)|\leq 2\frac{1-\tau_{\pi}}{\tau_{\pi}}M$. By Bernstein’s inequality for bounded distributions (e.g., Theorem 2.8.4 in Vershynin,, 2018) and the fact that $1/(a+b)\geq\min\{1/a,1/b\}/2$ for any $a,b>0$,
${\mathrm{pr}}(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta)$
$={E}\left[{\mathrm{pr}}(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta\mid\hat{Q}_{k,a})\right]$
$\leq 2{E}\left[\exp\left\{-\frac{n\eta^{2}}{2|I_{k}|P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}+\frac{4}{3}\frac{1-\tau_{\pi}}{\tau_{\pi}}Mn^{1/2}\eta}\right\}\right]$
$\leq 2{E}\left[\exp\left\{-\frac{\tau_{\pi}}{1-\tau_{\pi}}\frac{n\eta^{2}}{2|I_{k}|\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}+4Mn^{1/2}\eta/3}\right\}\right]$
$\leq 2{E}\left[\exp\left\{-\frac{\tau_{\pi}}{1-\tau_{\pi}}\left(\frac{n\eta^{2}}{4|I_{k}|\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\wedge\frac{3\sqrt{n}\eta}{8M}\right)\right\}\right]$
$=2{E}\left[\exp\{-\hat{\eta}(\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2})\}\right].$ (S6)
@@
$2{E}\left[\exp\{-\hat{\eta}(\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2})\}{\mathbbm{1}}(\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}\leq t)\right]$
$+2{E}\left[\exp\{-\hat{\eta}(\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2})\}{\mathbbm{1}}(\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}>t)\right]$
$\leq 2\exp\{-\hat{\eta}(t)\}+2{\mathrm{pr}}(\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}>t)$
$\leq 2\exp\{-\hat{\eta}(t)\}+2\hat{a}_{1}n^{\hat{d}_{1}}\exp\left\{-\hat{c}_{1}\left(\frac{t}{\hat{r}_{1}(n)}\right)^{\hat{q}_{1}}\right\}+2\hat{b}_{1}\sqrt{\log n/n}.$
Taking $t=\hat{r}_{1}(n)[(q+\hat{d}_{1})\hat{c}_{1}^{-1}\log n]^{1/\hat{q}_{1}}$ so that $\eta^{\prime}=\hat{\eta}(t)$ yields the desired inequality.
@@
Let $k\in\{1,\ldots,K\}$ be arbitrary. It holds that ${E}[\hat{\psi}_{k,a}]={E}[\hat{\psi}_{a}]=\psi_{*,a}$,
${\mathrm{var}}(\hat{\psi}_{k,a})$ $\leq\frac{1}{|I_{k}|}\left\{\sigma_{\#,a}+\sqrt{{E}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}\right]}\right\}^{2},$
${\mathrm{var}}(\hat{\psi}_{a})$ $\leq\left\{\frac{\sigma_{\#,a}}{\sqrt{n}}+\sum_{k=1}^{K}\frac{|I_{k}|}{n}\sqrt{{E}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\frac{(\hat{Q}_{k,a}-Q_{\#,a})^{2}}{|I_{k}|}\right]}\right\}^{2}.$
@@
${\mathrm{var}}(\hat{\psi}_{k,a})$ $\leq\frac{1}{|I_{k}|}\left\{\sigma_{\#,a}+\sqrt{\frac{1-\tau_{\pi}}{\tau_{\pi}}{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\right\}^{2},$
${\mathrm{var}}(\hat{\psi}_{a})$ $\leq\left\{\frac{\sigma_{\#,a}}{\sqrt{n}}+\sum_{k=1}^{K}\frac{\sqrt{|I_{k}|}}{n}\sqrt{\frac{1-\tau_{\pi}}{\tau_{\pi}}{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\right\}^{2}.$
@@
Since $\hat{Q}_{k,a}$ and $P_{n,k}$ are based on independent data, by Lemma S5, ${E}[(P_{n,k}-P_{*})H_{a}(\hat{Q}_{k,a})]=0$, and so ${E}[\hat{\psi}_{k,a}]={E}[\hat{\psi}_{a}]=\psi_{*,a}$.
@@
${\mathrm{var}}((P_{n,k}-P_{*})H_{a}(\hat{Q}_{k,a}))={E}[P_{*}H_{a}(\hat{Q}_{k,a})^{2}]/|I_{k}|=\frac{1}{|I_{k}|}{E}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}\right].$
@@
${\mathrm{var}}(\hat{\psi}_{k,a})={E}[(\hat{\psi}_{k,a}-\psi_{*,a})^{2}]$ $\leq\left\{\sqrt{{\mathrm{var}}[(P_{n,k}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})]}+\sqrt{{\mathrm{var}}[(P_{n,k}-P_{*})H_{a}(\hat{Q}_{k,a})]}\right\}^{2}$
$=\frac{1}{|I_{k}|}\left\{\sigma_{\#,a}+\sqrt{{E}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}\right]}\right\}^{2},$
${\mathrm{var}}(\hat{\psi}_{a})={E}[(\hat{\psi}_{a}-\psi_{*,a})^{2}]$ $\leq\left\{\sqrt{{\mathrm{var}}((P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a}))}+\sum_{k=1}^{K}\sqrt{{\mathrm{var}}\left(\frac{|I_{k}|}{n}(P_{n,k}-P_{*})H_{a}(\hat{Q}_{k,a})\right)}\right\}^{2}$
$=\left\{\frac{\sigma_{\#,a}}{\sqrt{n}}+\sum_{k=1}^{K}\frac{|I_{k}|}{n}\sqrt{{E}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\frac{(\hat{Q}_{k,a}-Q_{\#,a})^{2}}{|I_{k}|}\right]}\right\}^{2}.$
@@
${\mathrm{pr}}\left(|\hat{\sigma}_{k,a}^{2}-{E}[\hat{\sigma}_{k,a}^{2}]|>\frac{n}{|I_{k}|}\sum_{j=1}^{3}\lambda_{j}\right)$
@@
$\bar{\hat{\mathcal{R}}}(\lambda_{2}):=16M^{2}\frac{|I_{k}|^{2}}{n^{2}\lambda_{2}^{2}}\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}.$
Under Condition 8, $\bar{\hat{\mathcal{R}}}(\lambda_{2})$ may be replaced by
$\bar{\hat{\mathcal{R}}}^{\prime}(\lambda_{2}):=\hat{a}_{2}n^{\hat{d}_{2}}\exp\left\{-\hat{c}_{2}\left(\frac{n\lambda_{2}}{|I_{k}|\hat{r}_{2}(n)}\right)^{\hat{q}_{2}}\right\}+\hat{b}_{2}\sqrt{\log n/n}.$